OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now? EXPLAI IT OUT

OpenStudy (anonymous):

Let Tim's present age = x years Then Coach's age = x + 10 3 years ago Tim's age = (x - 3) years and Coach's age = (x + 10 -3) = (x + 7) years According to the question (x + 7) = 2*(x - 3) x + 7 = 2x - 6 7 + 6 = 2x - x 13 = x So Tim's present age = 13 years